# Batch Predictive Inference

### v4

@@ 4.1 Simultaneous predictive inference of multiple unobserved outcomes
We generate the data according to the distribution $X\sim N_{p}(\mu_{x},5\cdot I_{p}),Y\mid X\sim\mathcal{N}(\beta_{1}^{\top}X+(\beta_{2}^{\top}X)^{2},|\beta_{3}^{\top}X|^{2})$, where we set the dimension as $p=20$, and the mean vectors $\mu_{x}$ and $\beta_{1},\beta_{2},\beta_{3}$ are randomly generated by drawing each component from uniform distributions over the unit interval. First, we generate a training dataset of size $n_{\text{train}}=200$, and then fit a random forest regression estimator to estimate the mean function $\hat{\mu}(\cdot)$.
@@ 4.2 Selection with error control
Next, we illustrate the performance of batch PI procedure for the selection task described in 3.2. We generate the data from the distribution $X\sim N_{p}(\mu_{x},5\cdot I_{p}),Y=\log(1+\exp(\beta^{\top}X+\sigma Z)),\text{ where }Z\sim\mathcal{N}(0,1)$. The dimension is set to $p=20$, $\sigma=3$, and the mean vectors $\mu_{x}$ and $\beta$ are generated by drawing each component from uniform distributions over the unit interval. We consider the task of selecting individuals with $Y>5$, while controlling the number of false claims, i.e., the number of individuals selected whose actual outcome is five or less.
We first generate a training data of size $n_{\text{train}}=500$, and then fit a random forest regression to construct the score function $s:(x,y)\mapsto\hat{\mu}(x){\mathbbm{1}}\left\{{y\leqslant 5}\right\}$. Next, we repeat the process of generating calibration data of size $n=1000$ and test data of size $m=100$, 500 times. In each trial, we run the selection procedure (24) at level $\alpha=0.1$ and $0.2$, with $\eta=0,2,4,6,8,10$. We record the number of false claims, as well as the number of true claims in each trial. The results are summarized in Table 2 and Figure 3, illustrating that the proposed procedure controls the number of false claims across various target levels $\eta$, satisfying the guarantee (25).
$\eta=0$ $\eta=2$ $\eta=4$ $\eta=6$ $\eta=8$ $\eta=10$
[…]
{# false $\mathbb{P}\left\{{\text{\# false claims}>\eta}\right\}$ 0.084 (0.0006) 0.092 (0.0006) 0.104 (0.0006) 0.102 (0.0006) 0.122 (0.0007) 0.092 (0.0006)
[…]
{# false $\mathbb{P}\left\{{\text{\# false claims}>\eta}\right\}$ 0.172 (0.0008) 0.204 (0.0008) 0.212 (0.0008) 0.206 (0.0008) 0.206 (0.0008) 0.194 (0.0008)
Table 2: Mean of the number of false claims, probability of the number of false claims being larger than the target level $\eta$, and the power of the batch PI-based selection procedure, with standard errors, for $\eta=0,2,4,6,8,10$ and $\alpha=0.1$.
Figure 3: Number of false claims, probability of the number of false claims being larger than the target level $\eta$, and the power of the batch PI-based selection procedure, under $\eta=0,2,4,6,8,10$ and $\alpha=0.1,0.2$.
@@ 4.3 Inference on counterfactual variables
    Simultaneous inference on quartiles: Find $L,U$ such that $\mathbb{P}\left\{{L\leqslant Y_{(10)}^{a=0}\text{ and }Y_{(30)}^{a=0}\leqslant U}\right\}\geqslant 1-\alpha$.
Here, $Y_{(\zeta)}^{a=0}$ denotes the $\zeta$-th smallest value among $\{Y_{n+j}^{a=0}:j=1,2,\cdots,m\}$.
@@
We repeat the process of generating the calibration and test sets, and then applying the procedures at levels $\alpha=0.05,0.075,0.1,\cdots,0.2$, 500 times; and then compute the coverage rates. The results are summarized in Table 3 and 4, illustrating that the procedure tightly attains the target coverage rate.
Median 0.970 (0.0076) 0.940 (0.0106) 0.924 (0.0119) 0.906 (0.0131) 0.878 (0.0147) 0.852 (0.0159) 0.834 (0.0167)
Quartiles 0.972 (0.0074) 0.954 (0.0094) 0.932 (0.0113) 0.904 (0.0132) 0.868 (0.0152) 0.840 (0.0164) 0.814 (0.0174)
Table 3: Coverage rates of the batch PI prediction sets for counterfactual quartiles (upper: median, lower: quartiles) at different levels, with standard errors.
@@
Figure 4: Coverage rates of the batch PI prediction sets for the median and the quartiles of counterfactual variables at different levels. The dotted line corresponds to $y=x$ line.
@@
Next, we investigate the task of inference on the mean of counterfactual variables, where we aim to construct a bound $B$ that satisfies $\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}Y_{n+j}^{a=0}\leqslant B}\right\}\geqslant 1-\alpha$. We perform the experiment with a calibration size of $n=100$ and the test sizes of $m=5$ and $m=10$. The calibration size after rejection sampling is smaller—around 40 in this experiment. Thus, neither the naive method (which requires a sufficiently large calibration-to-test ratio) nor the concentration-based method (which requires large calibration and test sizes) is useful—they both provide trivial prediction sets of $[0,1]$.
@@
We repeatedly generate the data and run the batch PI procedure with the dynamic programming approach 4 500 times (which uses the rank-ordering function $\tilde{h}(r_{1},\cdots,r_{m})=\sum_{j=1}^{m}r_{j}$), and compute the coverage rate. The results are shown in Table 4 and Figure 5.
$m=5$ 0.976 (0.0069) 0.966 (0.0081) 0.954 (0.0094) 0.938 (0.0108) 0.928 (0.0116) 0.926 (0.0117) 0.916 (0.0124)
$m=10$ 0.982 (0.0060) 0.972 (0.0074) 0.962 (0.0086) 0.954 (0.0094) 0.848 (0.0099) 0.840 (0.0106) 0.824 (0.0119)
Table 4: Coverage rates of the prediction set for the mean of counterfactual variables at different levels, with standard errors.
Figure 5: Coverage rates of the prediction set for the mean of counterfactual variables for test sizes five and ten, at different levels. The dotted line corresponds to $y=x$ line.
The results show that the batch PI prediction set achieves the coverage guarantee, though it is a bit conservative. This conservativeness is partly due to the discrepancy between the the rank-ordering function $\tilde{h}$ and the actual ordering of the $h$ values, which we do not have access to in practice.
@@
To further illustrate this, we empirically examine the coverage rates of the batch PI prediction sets for the mean of the test scores under various score distributions with bounded support, with calibration and test sizes set to $n=40$ and $m=10$, respectively. Figure 6 demonstrates that the batch PI procedure achieves the target coverage guarantee across different distributions, though with varying levels of tightness. While the prediction set is designed to ensure a distribution-free guarantee—controlling for worst-case scenarios—it may be conservative in general. Nonetheless, these prediction sets remain useful in some sense, as neither naive methods nor concentration-based methods provide nontrivial prediction sets in this setting.
@@
In Appendix D, we provide simulation results in the setting where we do not have access to the true propensity score and instead use an estimate of the propensity score in the procedure. These results demonstrate that our methodology remains robust, yielding similar results even when relying on the estimates.
@@ 5 Empirical data illustration
Next, we illustrate the performance of the batch PI procedure by applying it to a drug-target interaction (DTI) dataset. We use the dataset and the pre-trained model from the DeepPurpose library (Huang et al., 2020). The original dataset has 16,486 observations in both the calibration and the test sets. The covariates consist of a pair of drug and target protein, and the response variable is the affinity score, which is a real-valued measure of the interaction between the drug and the target protein.
We first consider the task of constructing prediction sets for each unobserved outcome variable—as discussed in Section 3.1. To illustrate performance under moderate sample sizes, we create a calibration set of size 500 randomly drawn from the original calibration data. We then construct 160 test sets, each of size 100, using a total of 16,000 observations from the test set. Denoting the pretrained estimator by $\hat{\mu}$, we run the batch PI-based procedure (23) with the score $s:(x,y)\mapsto|y-\hat{\mu}(x)|$ at levels $\delta=0.1$ and $\alpha=0.05,0.1,0.15,0.2,0.25,0.3$. For comparison, we also run split conformal prediction at level $\delta=0.1$ for each of the test points. We compute the proportion of test sets (out of 160 total sets) where the coverage rate exceeds 0.9, as well as the mean coverage rate. The results are summarized in Table 5 and Figure 7.
batch PI ($\delta=0.1$) split conformal $(\text{level}=0.1)$
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.15$ $\alpha=0.2$ $\alpha=0.25$ $\alpha=0.3$
$\mathbb{P}\left\{{\textnormal{coverage}\geqslant 0.9}\right\}$ 0.9500 (0.0173) 0.9312 (0.0201) 0.8438 (0.0288) 0.7938 (0.0321) 0.7812 (0.0328) 0.7500 (0.0343) 0.6062 (0.0387)
$\mathbb{E}\left[{\textnormal{coverage}}\right]$ 0.9406 (0.0019) 0.9350 (0.0021) 0.9249 (0.0023) 0.9178 (0.0024) 0.9152 (0.0024) 0.9131 (0.0024) 0.9040 (0.0024)
Table 5: The proportion of test sets with test-coverage being larger than $0.9$, and the mean coverage rate of the batch PI-based procedure at levels $\alpha=0.05,0.1,\cdots,0.3$ and $\delta=0.1$, along with the split conformal-based procedure at level $0.1$, with standard errors.
@@
Next, we examine the task of selecting drug-target pairs with high scores, following the discussion in Section 3.2. We construct a calibration set of size 2000, and 160 test sets of size 100. We aim to select drug-protein pairs whose corresponding scores exceed seven, which roughly corresponds to the top 20% of the entire set. We run the procedure (24) at levels $\alpha=0.05,0.1,0.15,0.2,0.25,0.3$ and target numbers of false claims $\eta=0,3,5$ (recall that the procedure at $\eta=0$ controls a quantity analogous to the family-wise error rate (FWER)). The results are shown in Table 6 and Figure 8, illustrating that the batch PI procedure achieves the target guarantee at various levels.
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.15$ $\alpha=0.2$ $\alpha=0.25$ $\alpha=0.3$
$\eta=0$ 0.0437 (0.0162) 0.1250 (0.0262) 0.1938 (0.0313) 0.2125 (0.0324) 0.2250 (0.0331) 0.2375 (0.0337)
{# false $\mathbb{P}\left\{{\textnormal{\# false claims}\geqslant\eta}\right\}$ $\eta=3$ 0.0312 (0.0138) 0.0625 (0.0192) 0.1250 (0.0262) 0.1625 (0.0293) 0.2375 (0.0337) 0.2562 (0.0346)
$\eta=5$ 0.0125 (0.0088) 0.0625 (0.0192) 0.1375 (0.0273) 0.2188 (0.0328) 0.2688 (0.0352) 0.3000 (0.0363)
Table 6: The proportion of test sets whose number of false claims exceeds the target $\eta$, at levels $\alpha=0.05,0.1,\cdots,0.3$ and $\eta=0,3,5$, with standard errors.
Figure 8: The proportion of test sets whose number of false claims exceeds the target $\eta$, at levels $\alpha=0.05,0.1,\cdots,0.3$ and $\eta=0.3,5$. The dotted lines represent the $y=x$ line.
@@ 6 Discussion
Many open questions remain. For inference on one test point, several works have explored developing new distribution-free procedures that can achieve stronger targets or operate under more complex data structures. Examples include attaining training- or test-conditional coverage guarantees, or developing methods that work with non-exchangeable data. Similar questions can be asked for joint inference on multiple objects. For example, can we achieve batch-conditional inference, and what kind of conditional coverage can be controlled? If we have a hierarchical structure in the data involving groups of observations, how can we perform inference for new groups?
Another important direction would be to expand the range of targets that the inference framework can address. For example, if one is more interested in the proportion rather than the number of false claims in the selection task, can we construct a procedure with a guarantee of the form, $\mathbb{P}\left\{{\text{FDP}>\alpha}\right\}<\delta$? Can we generally target a functional of the empirical distribution of the unobserved outcomes? We leave these questions to future work.
Acknowledgement
@@
Appendix A Naive approaches
A.1 Partitioning the calibration data
A potential approach to achieve (1) is to partition the calibration data, to obtain multiple groups of observations that are exchangeable with the test set. Specifically, suppose $n=mq+r$ where $q$ is a non-negative integer and $10\leqslant r\leqslant m-$ 1. Let $\tilde{Z}_{k}=\{Z_{(k-1)m+1},Z_{(k-1)m+2},\ldots,Z_{km}\}$ for $k\in[q]$ and $\tilde{Z}_{\textnormal{test}}=\{Z_{n+1},\ldots,Z_{n+m}\}$. Then it is clear that $g(\tilde{Z}_{1}),g(\tilde{Z}_{2}),\ldots,g(\tilde{Z}_{q}),g(\tilde{Z}_{\textnormal{test}})$ are exchangeable, and thus we can apply split conformal prediction to obtain the following prediction set for $g(\tilde{Z}_{\textnormal{test}})$:
$\widehat{C}(\mathcal{D}_{n})=\left[Q_{\beta}^{\prime}\left(\sum_{k=1}^{q}\frac{1}{q+1}\delta_{g(\tilde{Z}_{k})}+\frac{1}{q+1}\delta_{\infty}\right),Q_{1-\gamma}\left(\sum_{k=1}^{q}\frac{1}{q+1}\delta_{g(\tilde{Z}_{k})}+\frac{1}{q+1}\delta_{\infty}\right)\right],$ (27)
A.2 Extending split conformal prediction
where $S_{i}=s(Z_{i})$. For $s_{1}\leqslant s_{2}\leqslant\ldots\leqslant s_{m}$, we define $h(s_{1},s_{2},\ldots,s_{m})$ as $\sup h$ if $s_{m}=\sup s$ and $h$ is not well-defined, e.g., $\sup s=+\infty$ and $h(s_{1},\ldots,s_{m})=\sum_{j=1}^{m}s_{i}$. Similarly, we define $h(s_{1},s_{2},\ldots,s_{m})$ as $\inf h$ if $s_{1}=\inf s$ and $h$ is not well-defined; while noting that only one of the two cases can occur below. Then, the naive split conformal prediction set $\widehat{C}(\mathcal{D}_{n})$ is defined as:
$\widehat{C}(\mathcal{D}_{n})=\left[Q_{\beta}^{\prime}\left(\sum_{1\leqslant i_{1}<\ldots<i_{m}\leqslant n+m}\frac{1}{\binom{n+m}{m}}\delta_{h(\underline{S}_{i_{1}},\ldots,\underline{S}_{i_{m}})}\right),Q_{1-\gamma}\left(\sum_{1\leqslant i_{1}<\ldots<i_{m}\leqslant n+m}\frac{1}{\binom{n+m}{m}}\delta_{h(\overline{S}_{i_{1}},\ldots,\overline{S}_{i_{m}})}\right)\right],$ (29)
However, this approach still faces limitations unless $n\gg m$. For instance, consider the scenario where $n=m$ and $\sup s=+\infty$. Then half of the $(\bar{S}_{i})_{1\leqslant i\leqslant n+m}$ values are +$\infty$+ ∞, likely leading to a trivial upper bound in (29).
A.3 Extending full conformal prediction
To avoid the issue of having a large mass at $\infty$ or $\infty$- ∞, one may try to construct a full conformal-type prediction set instead of relying on split conformal-type constructions. For example, we can first construct a joint prediction set for $(y_{(n+1):(n+m)})$ as
$\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})\\ =\left\{\tilde{y}=(y_{(n+1):(n+m)}):h(S_{(n+1):(n+m)}^{\tilde{y}})\leqslant Q_{1-\alpha}\left(\sum_{1\leqslant i_{1}<\ldots<i_{m}\leqslant n+m}\frac{1}{\binom{n+m}{m}}\delta_{h(S_{i_{1}}^{\tilde{y}},\ldots,S_{i_{m}}^{\tilde{y}})}\right)\right\},$ (30)
where $S_{i}^{\tilde{y}}=s^{\tilde{y}}(X_{i},Y_{i})$ and $s^{\tilde{y}}$ is the nonconformity score constructed from $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and $(X_{n+1},y_{n+1}),\ldots,(X_{n+m},y_{n+m})$—this step is essentially equivalent to the transductive conformal prediction Vovk [2013c]. Then the prediction set for $g(\{Z_{n+1},\ldots,Z_{n+m}\})$ can be constructed as $\widehat{C}(\mathcal{D}_{n})=\left\{g(\{(X_{n+1},y_{n+1}),\ldots,(X_{n+m},y_{n+m})\})):(y_{(n+1):(n+m)})\in\widehat{C}_{n}(X_{n+1},\ldots,X_{n+m})\right\}$.
A.4 Naive method: extending weighted conformal prediction

### v5

@@ 4.1 Simultaneous predictive inference of multiple unobserved outcomes
We generate the data according to the distribution
$X\sim N_{p}(\mu_{x},5\cdot I_{p}),Y\mid X\sim\mathcal{N}(\beta_{1}^{\top}X+(\beta_{2}^{\top}X)^{2},|\beta_{3}^{\top}X|^{2}),$
where we set the dimension as $p=20$, and the mean vectors $\mu_{x}$ and $\beta_{1},\beta_{2},\beta_{3}$ are randomly generated by drawing each component from uniform distributions over the unit interval. First, we generate a training dataset of size $n_{\text{train}}=200$, and then fit a random forest regression estimator to estimate the mean function $\hat{\mu}(\cdot)$.
@@
Next, we compare the batch PI-based method with the baseline Markov inequality-based method discussed in Remark 3, which attains the same guarantee. We follow the same steps of the previous simulation, while additionally applying the baseline method, at three different pairs of levels: $(\alpha,\delta)=(0.1,0.1)$, $(0.2,0.2)$, and $(0.3,0.3)$. Figure 3 shows the widths of the prediction sets from the two methods across different trials, illustrating that the batch PI-based method provides significantly shorter prediction intervals.
Figure 3: Prediction interval widths from the batch PI-based method and the conformal prediction with Markov-based level adjustment at different levels.
@@ 4.2 Selection with error control
Next, we illustrate the performance of batch PI procedure for the selection task described in 3.3. We generate the data from the distribution
$X\sim N_{p}(\mu_{x},5\cdot I_{p}),Y=\log(1+\exp(\beta^{\top}X+\sigma Z)),\text{ where }Z\sim\mathcal{N}(0,1).$
The dimension is set to $p=20$, $\sigma=3$, and the mean vectors $\mu_{x}$ and $\beta$ are generated by drawing each component from uniform distributions over the unit interval. We consider the task of selecting individuals with $Y>5$, while controlling the number of false claims, i.e., the number of individuals selected whose actual outcome is five or less.
We first generate a training data of size $n_{\text{train}}=500$, and then fit a random forest regression to construct the score function $s:(x,y)\mapsto\hat{\mu}(x){\mathbbm{1}}\left\{{y\leqslant 5}\right\}$. Next, we repeat the process of generating calibration data of size $n=1000$ and test data of size $m=100$, 500 times. In each trial, we run the selection procedure (20) at level $\alpha=0.1$ and $0.2$, with $\eta=0,2,4,6,8,10$. We record the number of false claims, as well as the number of true claims in each trial. The results are presented in Figure 4, illustrating that the proposed procedure controls the number of false claims across various target levels $\eta$, satisfying the guarantee (21).
Figure 4: Number of false claims, probability of the number of false claims being larger than the target level $\eta$, and the power of the batch PI-based selection procedure, for $\eta=0,2,4,6,8,10$ and $\alpha=0.1,0.2$.
Next, we compare the power of the proposed procedure and the methods based on Jin and Candès (2023b); Jin and Ren (2024), discussed in Section 3.3.1. We follow the same steps for the experiment as before but additionally run the procedures based on the conformal p-values (23) and (24), at levels $\alpha=0.05,0.075,0.1,\ldots,0.3$ and target false discovery bounds $\eta=0,5,10$. The results are shown in Figure 5, illustrating that the proposed method has significantly higher power than the conformal p-value-based methods in most settings.
Figure 5: Power of three procedures that control the $\eta$-FWER at different levels: (1) Batch PI-based procedure, (2) Procedure with conformal p-values by Jin&Candes (Jin and Candès, 2023b), (3) Procedure with conformal p-values by Jin&Ren (Jin and Ren, 2024).
@@ 4.3 Inference on counterfactual variables
    Inference on quartiles: Find $L,U$ such that $\mathbb{P}\left\{{L\leqslant Y_{(10)}^{a=0}\text{ and }Y_{(30)}^{a=0}\leqslant U}\right\}\geqslant 1-\alpha$.
Here, $Y_{(\zeta)}^{a=0}$ denotes the $\zeta$-th smallest value among $\{Y_{n+j}^{a=0}:j=1,2,\ldots,m\}$.
@@
We repeat the process of generating the calibration and test sets, and then applying the procedures 500 times, at levels $\alpha=0.025,0.05,0.075,\ldots,0.15$. Then we compute the coverage rates. For comparison, we also apply the baseline methods discussed in Section 2.1.1 (conformal+partitioning) and Section 2.1.3 (conformal+Bonferroni). The results are summarized in Figure 6. They show that our procedure tightly attains the target coverage rate—while the alternative methods output uninformative prediction sets.
@@
Figure 6: Coverage rates of the batch PI prediction sets for the median and the quartiles of counterfactual variables at different levels. The dotted line corresponds to $y=x$ line. Partitioning and the Bonferroni method both lead to trivial prediction sets that cover all possible outcomes, and have coverage equal to 100% (their lines overlap).
@@
Inference on the mean of counterfactuals. Next, we investigate the task of inference on the mean of counterfactual variables, where we aim to construct a bound $B$ that satisfies $\mathbb{P}\left\{{\frac{1}{m}\sum_{j=1}^{m}Y_{n+j}^{a=0}\leqslant B}\right\}\geqslant 1-\alpha$. We perform the experiment with a calibration size of $n=100$ and the test sizes of $m=5$ and $m=10$. The calibration size after rejection sampling is smaller—around 40 in this experiment. Thus, neither the partitioning-based method (which requires a sufficiently large calibration-to-test ratio) nor the concentration-based method (which requires large calibration and test sizes) is useful. For illustration, we also provide results for three baselines: conformal prediction with partitioning (see Section 2.1.1), conformal prediction with Bonferroni correction (see Section 2.1.3), and the concentration-based method (see Remark 2), and compare them with the batch PI-based procedure.
@@
We repeatedly generate the data and run the batch PI procedure with the dynamic programming approach from Section 2.3.2 (which uses the rank-ordering function $\tilde{h}(r_{1},\ldots,r_{m})=\sum_{j=1}^{m}r_{j}$) along with two comparison methods, for 500 trials. We then compute the resulting coverage rates. The results are shown in Figure 7.
Figure 7: Coverage rates of the prediction set for the mean of counterfactual variables obtained from batch PI and three baselines: conformal prediction with partitioning, conformal prediction with Bonferroni correction, and the concentration-based method, across different levels. The dotted line corresponds to $y=x$ line. Partitioning, Bonferroni and the concentration-based method all lead to trivial prediction sets that cover all possible outcomes, and have coverage equal to 100% (their lines overlap).
The results indicate that the batch PI prediction set satisfies the coverage guarantee, producing nontrivial prediction sets while the baseline method outputs nearly trivial prediction sets.
Understanding over-coverage. The coverage of our method is here higher than the nominal level. This reflects the inherent difficulty of the inference problem, rather than suggesting that the procedure is conservative. Observe that our inferential target is the following guarantee:
$\inf_{\text{all distributions }P}\mathbb{P}_{{(X_{i},Y_{i})_{i\in[n+m]}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P}}\left\{{\text{coverage event}}\right\}\geqslant 1-\alpha.$
The batch PI procedure aims to attain the above distribution-free guarantee by ensuring that the coverage rate exceeds $1-\alpha$ even in certain worst-case scenarios. As a result, in typical scenarios, the coverage may be higher than $1-\alpha$. For certain targets—e.g., inference on quantiles—Corollary 1 shows that we attain uniform tightness, i.e.,
$1-\alpha\leqslant\inf_{\text{all distributions }P}\mathbb{P}_{{(X_{i},Y_{i})_{i\in[n+m]}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P}}\left\{{\text{coverage event}}\right\}\\ \leqslant\sup_{\text{all distributions }P}\mathbb{P}_{{(X_{i},Y_{i})_{i\in[n+m]}\stackrel{{\scriptstyle\textnormal{i.i.d.}}}{{\sim}}P}}\left\{{\text{coverage event}}\right\}\leqslant 1-\alpha+O(\tfrac{1}{n}).$
However, for general targets, the tightness typically varies with the underlying distribution.
@@
To further illustrate this, we empirically examine the coverage rates of the batch PI prediction sets for the mean of the test scores under various score distributions with bounded support, with calibration and test sizes set to $n=40$ and $m=10$, respectively. Figure 8 demonstrates that the batch PI procedure achieves the target coverage guarantee across different distributions, though with varying levels of tightness. While the prediction set is designed to ensure a distribution-free guarantee—controlling for worst-case scenarios—it may be conservative in for particular data distributions. Nonetheless, these prediction sets remain useful and the only viable existing distribution-free method in this setting to our knowledge, as neither baseline methods nor concentration-based methods provide nontrivial prediction sets in this setting.
@@
In Section H, we provide simulation results in the setting where we do not have access to the true propensity score and instead use an estimate. These results demonstrate that our methodology similar results even when relying on the estimates.
@@ 5 Empirical data illustration
Next, we illustrate the performance of the batch PI procedure by applying it to a drug-target interaction (DTI) dataset to select high-scoring drug-target pairs. We use the dataset and the pre-trained model from the DeepPurpose library (Huang et al., 2020). The original dataset has 16,486 observations in both the calibration and the test sets. The covariates consist of a pair of drug and target protein, and the response variable is the affinity score, which is a real-valued measure of the interaction between the drug and the target protein.
We first consider the task of constructing prediction sets for each unobserved outcome variable—as discussed in Section 3.2. To illustrate performance under moderate sample sizes, we create a calibration set of size 500 randomly drawn from the original calibration data. We then construct 160 test sets, each of size 100, using a total of 16,000 observations from the test set. Denoting the pretrained estimator by $\hat{\mu}$, we run the batch PI-based procedure (19) with the score $s:(x,y)\mapsto|y-\hat{\mu}(x)|$ at levels $\delta=0.1$ and $\alpha=0.05,0.1,\ldots,0.3$. For comparison, we also run split conformal prediction at level $\delta=0.1$ for each of the test points. We compute the proportion of test sets (out of 160 total sets) where the coverage rate exceeds 0.9, as well as the mean coverage rate. The results are summarized in Figure 9.
@@
Next, we examine the task of selecting drug-target pairs with high scores, following the discussion in Section 3.3. We construct a calibration set of size 2000, and 160 test sets of size 100. We aim to select drug–protein pairs whose corresponding scores exceed a certain cutoff. We experiment with three cutoffs, chosen as the $q$-th quantiles of the score values in the training data—the remaining points after sampling 2000 points for the calibration set—with $q=0.7$, $0.8$, and $0.9$. [^15] This experimental design follows that of Jin and Candès (2023b). We run the procedure (20) at levels $\alpha=0.05,0.1,0.15,0.2,0.25,0.3$ and target numbers of false claims $\eta=0,3,5$ (recall that the procedure at $\eta=0$ controls a quantity analogous to the family-wise error rate (FWER)). The results are shown in Figure 10, illustrating that the batch PI procedure achieves the target guarantee at various levels.
Figure 10: The proportion of test sets whose number of false claims exceeds the target $\eta$, at levels $\alpha=0.05,\ldots,0.3$ and $\eta=0,3,5$, for three different cutoffs, with error bars. The dotted lines represent the $y=x$ line.
@@ 6 Discussion
Many open questions remain. For inference on one test point, several works have explored developing new distribution-free procedures that can achieve stronger targets or operate under more complex data structures. Examples include attaining training- or test-conditional coverage guarantees, or developing methods that work with non-exchangeable data. Similar questions can be asked for joint inference on multiple objects. For example, can we achieve batch-conditional inference, and what kind of conditional coverage can be controlled? If we have a hierarchical structure in the data involving groups of observations, how can we perform inference for new groups? We leave these questions to future work.
Acknowledgements
@@
Appendix A A simple example of our method and discussion of the choice of the rank ordering functions
Our method follows the logical progression outlined below:
  1. 1.
    For some random vector $R=(R_{1},\ldots,R_{m})\sim\text{Unif}(H)$, our target quantity is upper bounded by $h(S_{(R)}):=h(S_{(R_{1})},\ldots,S_{(R_{m})})$.
  2. 2.
    Once we construct a prediction set for $R$—i.e., a set $I\subset H$ such that $\mathbb{P}\left\{{R\in I}\right\}\geqslant 1-\alpha$, it follows that $\mathbb{P}\left\{{h(S_{(R)})\leqslant\max_{r\in I}h(S_{(r)})}\right\}\geqslant 1-\alpha$, and thus the desired coverage guarantee also holds.
For example, when $m=1$ and $h$ is the identity function so that our method reduces to standard conformal prediction, this corresponds to setting $I=\{1,2,\cdots,\lceil(1-\alpha)(n+1)\rceil\}$, and consequently, the upper bound for the test score is given by $\max_{r\in I}S_{(r)}=S_{(\lceil(1-\alpha)(n+1)\rceil)}$, which is exactly the bound provided by split conformal prediction. In this setting, the above $I$ is the one with the smallest $\max_{r\in I}h(S_{(r)})$, among all subsets of $H=[n+1]$ with coverage probability at least $1-\alpha$.
More generally, to obtain a short or tight prediction set, we want $\max_{r\in I}h(S_{(r)})$ to be small. To achieve this, the set $I$ should consist of those elements $r$ in $H$ whose corresponding $h(S_{(r)})$ values are small—which becomes a nontrivial task when $m>1$. For example, suppose $n=10$, $m=2$, and $h$ is the summation function, meaning that our target of inference is $S_{6}+S_{7}$. Then we consider the following $\binom{10+2}{2}=66$ sums:
$S_{(1)}+S_{(1)},S_{(1)}+S_{(2)},S_{(2)}+S_{(2)},\cdots,S_{(10)}+S_{(10)},S_{(10)}+\sup s,\sup s+\sup s.$ (26)
Mathematically, the set with the smallest $\max_{r\in I}h(S_{(r)})$ is the one containing the $\lceil 66\cdot(1-\alpha)\rceil$ smallest elements from the above list:
$I=\left\{1\leqslant r_{1}\leqslant r_{2}\leqslant n+1:S_{(r_{1})}+S_{(r_{2})}\leqslant Q_{1-\alpha}\left(\left\{S_{(r_{1}^{\prime})}+S_{(r_{2}^{\prime})}:1\leqslant r_{1}^{\prime}\leqslant r_{2}^{\prime}\leqslant n+1\right\}\right)\right\}.$ (27)
However, this choice does not yield a prediction set with valid coverage, since it essentially selects $I$ in a calibration-data-dependent manner, thereby breaking the logic in the above Step 2—specifically, the condition $\mathbb{P}\left\{{R\in I}\right\}\geqslant 1-\alpha$ does not hold for a data-dependent $I$.
Now, we require a set $I$ that does not depend on the data—for statistical validity—but still approximates the above ‘mathematically best’ $I$—to achieve a short and tight prediction set. The rank-ordering function $\tilde{h}$ was introduced to serve these two roles: it is independent of the data, but still tends to behave like $h$—since we want the resulting set $I$ to favor smaller elements in the list (26), so that $\max_{r\in I}h(S_{(r)})=\max_{(r_{1},r_{2})\in I}(S_{(r_{1})}+S_{(r_{2})})$ remains small. For example, in the paper, we discuss two strategies:
  1. 1.
    Rank-ordering functionally identical to the batch score: we use
    $I=\left\{1\leqslant r_{1}\leqslant r_{2}\leqslant n+1:r_{1}+r_{2}\leqslant Q_{1-\alpha}\left(\left\{r_{1}^{\prime}+r_{2}^{\prime}:1\leqslant r_{1}^{\prime}\leqslant r_{2}^{\prime}\leqslant n+1\right\}\right)\right\}.$
  2. 2.
    Rank ordering based on independent split: we use
    $I=\left\{1\leqslant r_{1}\leqslant r_{2}\leqslant n+1:\tilde{S}_{(r_{1})}+\tilde{S}_{(r_{2})}\leqslant Q_{1-\alpha}\left(\left\{\tilde{S}_{(r_{1}^{\prime})}+\tilde{S}_{(r_{2}^{\prime})}:1\leqslant r_{1}^{\prime}\leqslant r_{2}^{\prime}\leqslant n+1\right\}\right)\right\},$
    where $\tilde{S}_{i}$ are scores from an independent data split (of the same size).
In summary, the underlying intuition is to approximate the mathematically optimal—but statistically non-justified—prediction set $I$ (27) for the ranks, using the function $\tilde{h}$ that mimics $h$.
Appendix B Naive method: extending weighted conformal prediction